OpenStudy (anonymous):

The dimensions of a rectangular prism are the same as a Pythagorean triangle with legs of 7 and 24. The volume of the prism is

OpenStudy (anonymous):

since the area of the prism= area of the triangle hypotesues of triangle = 25 sin<B=24/25 <B=74 degrees. sinB/h=sin90/7 h=sinB*7/1 h=6.7 Area triangle= (1/2) base*height A=(1/2) 25*6.7=84 units square

OpenStudy (anonymous):

Okay! That makes more sense. Thankyou.

OpenStudy (anonymous):

ur welcome(:

OpenStudy (raden):

is there a diagram ?